OpenStudy (anonymous):

a pair of dice are tossed, determine the probability that at least one of the two dice shows a 1. Can someone please show me how to solve this.

OpenStudy (cwrw238):

since there are 6 sides to a dice the probability of throwing 1 is 1/6 and probability of not throwing a 1 is 5/6 when throwing 2 dice the possiblities for 'at least one 1' are (1) 1 and not 1 or (2) not 1 and 1 or (3) 1 and 1 probability for (1) = 1/6 * 5/6 = 5/36 (2) = 5/6 * 1/6 = 5/36 (3) = 1/6 * 1/6 = 1/36 adding required probability = 5/36 + 5/36 + 1/36 = 11/36

OpenStudy (cwrw238):

a tree diagram illustrates this: